Lynn has a starting salary of $30,000 per year. Her company offers her a raise of 5% each year.
a. Find a rule for Lynn's salary in her nth year. How much will she be earning in her fifth year?
b. Another company offers Lynn the same starting salary with a raise of $2,000 each year. Lynn plans to stay with the same company for 10 years. Which company has the better offer? Explain your answer.

Answer by richwmiller(17219)

Tn = t * r^(n - 1)
Tn = 30000 * 1.05^(n-1)

end of yr 1 31500
yr 2 33075
yr 3 34728.75
yr 4 36465.1875
yr 5 38288.446875
yr 10 48866.8388033232421875
a10 = 30000 + (9)*2000
a10 = 30000 + 18000
a10 = 48000